URGENT!
37 min left

Pre re oswer uget eecab y Clicron, Inc. CurreadyBrown uses 20 percent debt in is capital structure, but Cicron plans to increae the debt ratio to 25 percent if the scquistion is consummated The after-tar cost of debt capital for Brown is estimated to be 8percent, which holds constant under either apital structure. The cost of equity sfer the acquisition is expected to be 22 percent. The curent ket value of Brown oustanding debt is S75 million, ll f which will be assumed by Clcron. Cicron intends to pay S225 million in cash and common stock for all f Browns stock in sddition to suming al of Browas debt. Currendy the murket price of Browas common stock is $200 millio. Selected items from Browns financial data are as follows 2004 2005 2006 2007 THEREAFTER (MILLIONS) $260 $265 $280 $290 $300 Net sales dministrative and selling expenses 25 25 25 30 3 Depreciation Capital expendlitures 15 17 18 2330 22 8 18 20 22 Insddition, the cost of goods sold runs 50percent of sales andthe marginal tax rate is 34 percent. Compute the net present value of the acquisition Required: a. Compute NPV of the acquisition b. Make decision to acquire or not Explain
Answer #1

WACC for Brown if the acquisition goes through:

WACC = (weight(debt) * Cost(debt) * (1 - tax rate)) + (weight(equity) * cost(equity))

weight(debt) after acquisition = 25%

cost(debt) = 8%

tax rate = 34%

weight(equity) = 1 - weight(debt) = 1 - 0.25 = 0.75 = 75%

cost(equity) after acquisition = 22%

WACC = (0.25 * 0.08 * (1 - 0.34)) + (0.75 * 0.22)

= (0.25 * 0.08 * 0.66) + (0.75 * 0.22)

= 0.0132 + 0.165 = 0.1782

WACC = 17.82%

For computing PV of Brown:

(All values are in $ million, unless states otherwise)

Particulars 2004 2005 2006 2007 Thereafter Formula
Net Sales 260 265 280 290 300
(-) Cost of goods sold (COGS) (130) (132.5) (140) (145) (150) 50% of net sales
Gross profit (GP) 130 132.5 140 145 150 Net sales - COGS
(-) Administrative & Selling Expenses (25) (25) (25) (30) (30) Given in question
(-) Depreciation (Dep) (15) (17) (18) (23) (30) Given in question
Earnings before Interest & taxes (EBIT) 90 90.5 97 92 90 GP - Admin - Dep
(-) Taxes at 34% (30.6) (30.77) (32.98) (31.28) (30.6)
Net Income 59.4 59.73 64.02 60.72 59.4 EBIT - taxes
(+) Depreciation 15 17 18 23 30
Net Operating Profit after tax (NOPAT) 74.4 76.73 82.02 83.72 89.4 Net Income + Dep
(-) Capital Expenditure (Capex) (22) (18) (18) (20) (22) Given in question
Free cash flow to the firm (FCFF) 52.4 58.73 64.02 63.72 67.4 NOPAT - Capex
Discounting factor at WACC of 17.82%

1/(1 +0.17821)

= 0.8488

1/(1 +0.17822)

= 0.7204

1/(1 +0.17823)

= 0.6114

1/(1 +0.17824)

= 0.5189

1/(1 +0.17825)

= 0.4405

PV of FCFF 44.48 42.31 39.14 33.06 29.69 FCFF * Discounting factor

Sum of PV of FCFF = 44.48 + 42.31 + 39.14 + 33.06 + 29.69

= 188.68

NPV acquisition = PV of FCFF - PV of cash outflow by Cicron

= 188.68 - 225

= -36.32

Implying that the present values of potential future inflows from brown are values at $188.68 million which is lower than the market value of the common stock of brown at $200 million as well as amount of $225 million which Cicron is expected to pay for the acquisition. in this scenario, the overall acquisition will lead to net cash outflow for Cicron and hence, Cicron should not acquire Brown.